If N is nil ideal of a ring R and A is ideal of R such that A contained in N then prove that A and N/A are also nil ideal
1
Expert's answer
2021-07-20T10:07:32-0400

Let "p\in N/A". Then "p=q+A" where "q\in N". Since "N" is a nil ideal, "\exist k\in \mathbb{Z}" such that "q^k=0". But "0\in A". So "q^k\in A", which implies that


"p^k=(q+A)^k=q^k+A=0+A"

the zero element of "N\/A". Hence "N\/A" is a nil ideal.